OpenStudy (anonymous):

Find the product of (x + 3)(x − 3).

ganeshie8 (ganeshie8):

we do the same

ganeshie8 (ganeshie8):

use this :- \(\large (a+b)(a-b) = a^2 - b^2\)

ganeshie8 (ganeshie8):

\(\large (x + 3)(x − 3) = ?\)

OpenStudy (anonymous):

distribute x^2-3x+3x-9 do u know what to do now or u need further assistance?

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

np :) glad i could help :)

OpenStudy (irishboy123):